OpenStudy (lucaz):

indefinite integral of (x^2+1)/x^2, how do I do this?

hartnn (hartnn):

is it just \(\dfrac{x^2+1}{x^2 }\) ?

OpenStudy (lucaz):

yes

hartnn (hartnn):

then \(1+\dfrac{1}{x^2} = 1+x^{-2}\)

hartnn (hartnn):

separating out the numerator

OpenStudy (lucaz):

oh, ok thanks

hartnn (hartnn):

can you finish ?

OpenStudy (lucaz):

I think so

hartnn (hartnn):

if you want me to verify your answer, then tell me what u get :)

OpenStudy (lucaz):

thank you